# Decouple and Decompose: Scaling Resource Allocation with DeDe

### v2

@@ 3.2 Decompose
To this point, the original resource allocation problem has been transformed into alternating minimization of two augmented Lagrangians (original objective plus penalty terms) with respect to $x$ and $z$, respectively. Next, we take the $x$-minimization step as an example to describe how it can be decomposed ($z$-minimization is similar).
As $x$-minimization focuses on the variable $x$ only, we can ignore terms related to $z$ in the augmented Lagrangian, such as the penalty terms from demand constraints. The resulting objective consists of three resource-related terms:
  • sum of per-resource allocation cost $f_{i}(x_{i*})$ from Equation 5;
  • sum of penalties for violating each resource constraint $R_{i}x_{i*}=r_{i}$ from Equation 2;
  • sum of penalties for violating $x=z$ from Equation 6.
The first two items are a sum of functions on per-resource allocation $x_{i*}$, while the third can also be rewritten as a sum of penalties for violating each $x_{i*}=z_{i*}$ (equivalent to $x=z$).
After rearranging, the optimization can be expressed as:
$\min_{x\in\mathcal{X}}\sum_{i}F_{i}(x_{i*}),$
where $F_{i}(x_{i*})$ consists of allocation cost, resource constraint penalty, and penalty for violating $x_{i*}=z_{i*}$. As each variable $x_{i*}$ is mutually exclusive and independent, the above problem is equal to optimizing per-resource allocations in parallel:
$\min_{x_{i*}\in\mathcal{X}_{i*}}F_{i}(x_{i*}),\quad\forall i,$
thereby decomposing the allocation of $x$ with $n\times m$ variables into $n$ per-resource allocations of $x_{i*}$, each with $m$ variables. To solve them, DeDe uses an off-the-shelf solver (cvxpy in our implementation) to find (exact or approximate) solutions while making liberal use of all available CPU cores.
@@
NetHint [14], Gavel [44], Teal [63], OneWAN [32], BLASTSHIELD [31], NCFlow [1], Cerebro [66], DOTE [47], Soroush  [42], POP [43]
@@ 4 Generality and Limitations
In this section, we start with the generality of DeDe, and then discuss its limitations, including scenarios with limited parallelism and chances of suboptimal solutions.
@@ 4.1 Generality
DeDe can be applied to a wide range of resource allocation problems. Table 1 presents a summary of real-world resource allocation problems in the recent literature, mostly from OSDI, SOSP, NSDI, and SIGCOMM in the past four years. Next, we explain why DeDe is applicable.
  • Variables.  Inheriting from ADMM, DeDe inherently supports real-number (float) allocation variables. Empirical evidence[53] and theoretical proofs [56, 38, 60] on ADMM also indicate that DeDe can effectively handle boolean and integer variables by projecting real-number solutions to desired data types over iterations.
  • Objective.  DeDe requires a max or sum of convex utility functions over per-resource or per-demand allocations. Most objectives used in systems exhibit convexity, such as the weighted allocation sum [44, 63], the logarithmic resource utility [3], and the quadratic allocation cost [64, 59].
  • Constraints.  DeDe assumes the linearity of resource and demand constraints, which holds in all our surveyed problems, e.g., the sum of demands allocated to a resource must not exceed its capacity. The constraints in these problems are per resource or per demand as well. Even under a constraint on multiple resources or demands, DeDe can be relaxed to groups of resources or demands. DeDe bundles these groups together into per-resource-group or per-demand-group allocations during decomposition.
Unlike POP [43], DeDe works with any workload, not only “granular” ones. In POP, the demands in each decomposed problem can access only a subset of resources, assuming each demand only requests a small fraction of interchangeable resources such that a resource subset suffices. This granular assumption can be unrealistic in practice, due to uneven demand distributions [28, 57] or limited resources that each demand is allowed to utilize [58, 33]. In contrast, each decomposed problem in DeDe can still access all the resources or demands; thus, the granular assumption is no longer required, affirming DeDe’s generality to various workload types.
@@ 4.2 Limitations
While DeDe applies to a variety of resource allocation problems, it might not always be the most suitable solution for every allocation problem.
Limited parallelism.  DeDe achieves parallelism by decomposing the full problem into per-resource and per-demand allocations. Therefore, when we cannot fully decompose the problem, DeDe will experience limited parallelism, although this does not happen in our survey (Table 1).
  • Non-separable objectives.  When the objective function is not a sum or max over allocation qualities per resource or per demand, we must bundle per-resource or per-demand allocations whose utilities stick together during decomposition. As a result, the partially decomposed problem will achieve limited parallelism.
  • Non-separable constraints.  When a resource constraint restricts the allocation across multiple resources, we must bundle the allocations for these resources during decomposition, which leads to limited parallelism. This also applies to demand constraints.
Suboptimal solutions.  It is possible for DeDe not to attain an ideal solution when the allocation problem deviates from the standard ADMM formulation.
  • Non-convex problems.  ADMM can solve non-convex optimization problems [53, 62, 68] empirically, providing theoretical guarantees only under specific assumptions[56, 38, 60]. As a result, DeDe can output suboptimal solutions in certain non-convex cases.
  • Nonlinear constraints.  When a constraint is nonlinear, we replace it by adding an indicator function to the objective. The indicator function is infinity when the constraint is violated or otherwise zero. However, this may alter the convexity of the objective function, resulting in the above non-convex problems.
  • Higher allocation dimension.  The allocation of resources to demands is encoded in a two-dimensional matrix $x$. Correspondingly, we separate the problem into two ADMM blocks for resources and demands. Introducing more dimensions, such as a temporal dimension [36], leads to more ADMM blocks and risks convergence failure [13].
@@ 5 Case Studies
In this section, we will present case studies on how DeDe can be applied to three resource allocation problems: traffic engineering, cluster scheduling, and load balancing, as also studied in POP [43].
@@ 5.1 Traffic Engineering
The traffic engineering problems [24, 54, 43, 42, 63] allocate traffic demand between datacenter nodes onto network links in a WAN topology.
Maximize total flow.  The problem focuses on maximizing the total flow allocation.
  • Variables.  The flow allocation from datacenter node pair $(s,t)$ onto network link $(u,v)$ is denoted as $x_{(u,v)(s,t)}\geq 0$. In common path-based traffic engineering, flows between node pair $(s,t)$ are allocated over links along pre-configured paths $P_{(s,t)}$ only.
  • Objective.  For a node pair $(s,t)$, the inflow it incurs to node $v$ is computed as $\text{inflow}(v,(s,t))=\sum_{(u,v)\in P_{s,t}}x_{(u,v)(s,t)}$ and outflow as $\text{outflow}(v,(s,t))=\sum_{(v,u)\in P_{s,t}}x_{(v,u)(s,t)}$. The total flow allocated for the node pair $(s,t)$ can be represented as $\text{inflow}(t,(s,t))$, which denotes the total inflow toward the destination $t$. Therefore, the objective of maximizing total flows between all node pairs can be expressed as the sum:
@@
  • Resource constraints.  For each link $(u,v)$, the total flow allocation from all node pairs must not exceed its link capacity $c_{u,v}$:
@@
Min-max link utilization.  The problem of min-max link utilization aims to allocate demands fairly over network links. The problem shares the same variables and linear demand constraints as the max total flow problem.
  • Objective.  The utilization of a link $(u,v)$ is the ratio of the total flow assigned to the link to its capacity. Our objective is to minimize the max link utilization across the network:
@@
Using DeDe.  The traffic engineering problem aligns the separable structure for DeDe: the objectives are the sum of allocation quality per demand or the maximum of allocation quality per link, respectively. Linear constraints are applied to each demand and each link. Therefore, DeDe can decompose the problem into $|E|$ per-link allocations and $|V|^{2}$ per-demand allocations for parallelism.
5.2 Cluster Scheduling
Cluster scheduling problems [61, 44, 43, 42] involves allocating jobs to computing resources (e.g., CPUs, GPUs) in heterogeneous clusters.
Max-min allocation.  This problem aims to distribute the heterogeneous cluster resources fairly among jobs based on their effective throughput.
  • Variables.  With jobs time-sliced onto the available heterogeneous computing resources [44, 43], the allocation is a matrix $x\in[0,1]^{n\times m}$ where $x_{ij}$ is the fraction of time each job $j$ spent on each computing resource type $i$.
  • Objective.  Given an allocation matrix $x$, we adopt a normalized effective throughput of job $j$ from POP [41], which is a linear function on $x_{*j}$. The objective is to maximize the minimum allocation of normalized effective throughput:
  • Resource constraints.  For each computing resource type $i$, the total time allocation cannot exceed the number of that computing resource type:
    $\sum_{j}x_{ij}z_{j}\leq\text{num\_workers}_{i},\quad\forall i,$
    where $z_{j}$ is the computing resources requested by job $j$.
  • Demand constraints.  For each job $j$, The sum of all fractions of time must be within 1:
    $\sum_{i}x_{ij}\leq 1,\quad\forall j.$
Proportional fairness.  The proportional fairness problem maximizes overall resource utilization while ensuring minimum service for each job. The problem shares the same variables, resource constraints, and capacity constraints.
Using DeDe.  The cluster scheduling exhibits the separable structure required by DeDe: The objective function is a minimum of allocation quality per job and a sum of allocation quality per job, respectively. Additionally, linear constraints are applied to each job and each computing resource. Thus, DeDe can decompose the problem into $n$ per-job allocations and $m$ per-demand allocations.
@@ 5.3 Load Balancing
Load balancing problems [16, 48, 52, 43] allocate data shards among servers to scale out query loads in a distributed store.
Minimize shard movements.  This problem aims to minimize shard movements across servers during load changes while keeping the load nearly balanced on each server.
  • Variables.  The allocation is a matrix $x\in[0,1]^{n\times m}$, where $x_{ij}$ is the fraction of data shard $j$ allocated to server $i$. Additionally, a binary matrix $x^{\prime}$ captures shard placement, where $x^{\prime}_{ij}=1$ if the data shard $j$ is on the server $i$, i.e., $x_{ij}>0$; otherwise $x_{ij}=0$.
  • Objective.  The movement of data shard $j$ allocated to server $i$ can be represent as $(1-T_{ij})x^{\prime}_{ij}f_{i}$, where $T_{ij}$ denotes the initial shard placement and $f_{j}$ denotes the memory footprint of data shard $j$. The objective of minimizing total shard movement can be expressed as the sum:
@@
  • Resource constraints.  For each server $i$, the total query load must be close to the average query load $L$; The total memory usage must be within memory capacity:
@@
  • Demand constraints.  Each data shard $j$ must allocate all the query load among servers:
@@
Using DeDe.  Given the separable structure, where the objective is the sum of allocation quality per data shard and linear constraints apply to each server, and each shard, DeDe can decompose the problem into $n$ per-server allocations and $m$ per-shard allocations.
@@ 6 Implementation of DeDe
Listing 1 showcases a simple resource allocation problem solved using our DeDe package. In line 4, we initially create a $N\times M$ non-negative allocation variable x, where x[i,j] represents the allocation of resource i to demand j. We then create a constraint for each resource in line 7–8 and a constraint for each demand in line 9–10, as well as an objective of minimizing the sum of x in line 13. Finally, we construct the problem in line 16–17 and solve it in line 18.
1import dede
2
3# Create allocation variables
4x = dede.Variable((N, M), nonneg=True)
5
6# Create the constraints
7resource_constrs = [
8 x[i,:].sum() <= 1 for i in range(N)]
9demand_constrs = [
@@
11
12# Create an objective
13obj = dede.Minimize(x.sum())
14
15# Construct and solve the problem
16prob = dede.Problem(
@@
18prob.solve(num_cpus=64)
@@
DeDe borrows the interface from cvxpy [17, 4]. DeDe inherits most of cvxpy’s methods, e.g. Variable(.) in line 4, Minimize(.) in line 13. One difference from cvxpy is that users must specify resource constraints and demands constraints separately (line 16–17) when initializing a problem. As a parallel solution, DeDe allows users to specify the number of CPU cores (line 18); if not specified, DeDe uses all available CPU cores.
Behind DeDe’s interface, the resource allocation problem is solved in the following three steps.
  • Problem parsing.  Inequality constraints are converted into equality constraints with a non-negative slack variable s., e.g., inequality x[0,:].sum() <= 1 can be converted to equality x[0,:].sum() + s == 1. The slack variable can be regarded as an additional part of the problem variables and be optimized together.
  • Problem building.  Resource constraints are organized into independent per-resource subgroups without shared variables, while demand constraints are similarly organized into separate independent per-demand groups. For each subgroup, a per-resource or per-demand problem is built using cvxpy, with the new objective formulated based on the augmented Lagrangian method outlined in § 3.2.
  • Problem solving.  Per-resource or per-demand problems in cvxpy are solved in parallel across multiple cores. Between ADMM iterations, only the parameters are updated, avoiding the need to rebuild the entire cvxpy problem. Similarly, for the same problem with varying resources and demands, only the corresponding parameters are updated.
We use Ray [40] to parallelize our computation across multiple cores. Python has limited parallelism in its multithreading implementation due to the notorious global interpreter lock (GIL). Instead, we choose Ray, which provides a wrapper for multiple Python interpreter processes, so as to avoid GIL. We also benefit from Ray’s production-grade implementation of inter-process coordination mechanisms to avoid building them from scratch using Python’s built-in low-level multiprocessing library.

### v3

@@ 3.2 Decompose
Up to this point, we have reformulated the original resource allocation problem into alternating minimization steps of the augmented Lagrangian with respect to $x$ and $z$. Next, we take the $x$-minimization step as an example to describe how the optimization is decomposed ($z$-minimization is similar).
@@
NetHint[12], Gavel[45], Teal[65], ONEWAN[33], BLASTSHIELD[32], NCFlow[1], Cerebro[68], DOTE[48], POP[44]
@@
Because the $x$-minimization step solves for $x$ only, we can ignore all terms in the augmented Lagrangian without $x$ (such as demand constraints). Thus, we rewrite Equation 6 as
$+\frac{\rho}{2}\big{\|}x-z^{(k)}+\lambda^{(k)}\big{\|}_{F}^{2}$
$=$ $\operatorname*{arg\,min}_{x\in\mathcal{X}}\,\sum_{i}\Bigl{(}f_{i}(x_{i*})+\frac{\rho}{2}\big{\|}R_{i}x_{i*}-r_{i}+\alpha_{i}^{(k)}\big{\|}_{2}^{2}$
Given that the new objective is a sum of disjoint terms across index $i$, its minimization can be decomposed into independent, per-resource subproblems over each $x_{i*}$:
$\begin{split}x_{i*}^{(k+1)}=&\operatorname*{arg\,min}_{x_{i*}\,\in\mathcal{X}_{i*}}\,\Bigl{(}f_{i}(x_{i*})+\frac{\rho}{2}\big{\|}R_{i}x_{i*}-r_{i}+\alpha_{i}^{(k)}\big{\|}_{2}^{2}\\ &+\frac{\rho}{2}\big{\|}x_{i*}-z_{i*}^{(k)}+\lambda_{i*}^{(k)}\big{\|}_{2}^{2}\Bigr{)},\quad\forall i.\end{split}$ (8)
This decomposition is made possible by the separable structure: (1) the original objective is separable, with a utility $f_{i}(x_{i*})$ defined for allocating each resource $i$, and (2) the resource constraints are likewise separable, as each constraint $R_{i}x_{i*}=r_{i}$ pertains only to the corresponding resource $i$.
As a result, the original optimization problem over $n\times m$ variables is broken down into $n$ independent subproblems, each involving only $m$ variables. DeDe solves each subproblem using off-the-shelf solvers, which return either exact or approximate solutions depending on the nature of $f$ (convex or not), whether $\mathcal{X}$ contains integers, etc. Since this decomposition allows up to $n$ parallel computations—often numbering in the thousands in large-scale settings—DeDe is capable of delivering substantial speedups assuming sufficient CPUs.
For linear programs, the original problem has a worst-case time complexity of approximately $O((n\cdot m)^{2.373})$ [13, 44]. After decomposition, DeDe processes $n$ subproblems, each with a complexity of $O(m^{2.373})$, reducing the overall complexity to $O(n\cdot m^{2.373})$. Nonetheless, this theoretical improvement assumes that ADMM converges within a constant number of iterations. In practice, the number of iterations depends on the choice of the penalty parameter $\rho$ (which acts like a “learning rate”) and other factors that influence convergence.
Similarly, the $z$-minimization step decomposes into independent, per-demand allocation subproblems over each $z_{*j}$:
$\begin{split}z_{*j}^{(k+1)}=&\operatorname*{arg\,min}_{z_{*j}}\,\Bigl{(}g_{j}(z_{*j})+\frac{\rho}{2}\big{\|}D_{j}z_{*j}-d_{j}+\beta_{j}^{(k)}\big{\|}_{2}^{2}\\ &+\frac{\rho}{2}\big{\|}x^{(k+1)}_{*j}-z_{*j}+\lambda_{*j}^{(k)}\big{\|}_{2}^{2}\Bigr{)},\quad\forall j.\end{split}$ (9)
@@ 4 Generality and Limitations
In this section, we discuss the generality of DeDe (conditions under which it applies) and its limitations (scenarios that may result in limited parallelism or suboptimal solutions).
@@ 4.1 Generality
DeDe is applicable to a wide range of real-world resource allocation problems. Table 1 summarizes these problems that appear in recent literature such as OSDI, SOSP, NSDI, and SIGCOMM. Below, we explain why DeDe is applicable.
  • Variables.  As an ADMM-based framework, DeDe natively supports continuous (real-valued) allocation variables. Empirical evidence[54] and theoretical proofs [57, 39, 61] on ADMM further indicate that DeDe should effectively handle boolean and integer variables by projecting real-valued solutions onto the appropriate domains during iterations.
  • Objective.  Most objectives in real systems are defined as the sum, max, or min of utility functions over the individual allocation of resources and demands. They also exhibit convexity. Examples include weighted sums [45, 65], logarithmic utilities [2], and quadratic costs [66, 60].
  • Constraints.  DeDe assumes linear constraints on resources and demands—a condition that holds across all the surveyed use cases. These constraints are commonly expressed at the per-resource or per-demand level. Even for constraints that do span multiple resources or demands, DeDe can turn them into disjoint constraint groups and apply decomposition at the group level (at the cost of reduced parallelism).
Unlike POP [44], DeDe works with any workload, not only “granular” ones. In POP, the demands in each subproblem can only access a subset of resources, assuming each demand only requests a small fraction of interchangeable resources such that a resource subset suffices. This granular assumption can be unrealistic in practice, due to uneven demand distributions [29, 58] or specific resources that each demand is allowed to utilize [59, 34]. In contrast, each subproblem in DeDe can still access all the resources or demands, resulting in generalization to various workload distributions.
@@ 4.2 Limitations
While DeDe applies to a variety of resource allocation tasks, it might not always be the optimal approach for all scenarios.
Limited parallelism.  DeDe’s parallelism relies on its ability to decompose the problem into independent subproblems. When such decomposition is not fully achievable, the degree of parallelism diminishes. Although this limitation does not manifest in the set of real-world problems we surveyed, we come up with some examples below.
  • Non-separable objectives.  An objective is non-separable when utility depends on the interaction between different resources or demands, e.g., if a job attains higher throughput when it has exclusive access to a GPU compared with when sharing it with other jobs.
  • Non-separable constraints.  A constraint is non-separable when it spans multiple resources and multiple demands simultaneously. For example, this occurs when the jobs submitted by each user are collectively limited by a user-specific GPU hour quota. To restore decomposability, DeDe has to treat all jobs (the original demands) from the same user as a single aggregated demand. However, this aggregation reduces the granularity of parallelism, from the job level to the user level.
Suboptimal solutions.  DeDe may fail to reach the optimal solution when the allocation problem deviates from the assumptions underlying standard ADMM.
  • Non-convex problems.  Although ADMM has demonstrated empirical success on various non-convex problems [54, 64, 70], its theoretical convergence is guaranteed only under specific conditions [57, 39, 61]. Consequently, DeDe may converge to suboptimal solutions in non-convex settings.
  • Nonlinear constraints.  When a constraint is nonlinear, DeDe has to reformulate it with an indicator function in the augmented Lagrangian. The indicator function returns zero if the constraint is satisfied and infinity otherwise. However, this approach destroys the convexity of the objective, resulting in the aforementioned non-convex problems.
  • Higher allocation dimension.  DeDe assumes a two-dimensional allocation matrix that maps resources to demands, leading to a two-block ADMM formulation. Introducing additional dimensions such as a temporal dimension [37] will increase the number of variable blocks in ADMM and thus might lead to convergence issues [11].
@@ 5 Case Studies
In this section, we present case studies of DeDe applied to three representative resource allocation problems [44]: cluster scheduling, traffic engineering, and load balancing.
5.1 Cluster Scheduling
Cluster scheduling problems [63, 45, 44, 43] involve allocating computing jobs to different resource types in heterogeneous clusters. The job scheduling example introduced earlier is a simplified version of this problem.
Max-min allocation.  This variant aims to distribute the heterogeneous cluster resources fairly among jobs such that the minimum throughput is maximized.
  • Variables.  Jobs can be time-sliced across available resource types [45, 44]. Thus, the allocation plan is a matrix $x\in[0,1]^{n\times m}$, where $x_{ij}$ is the fraction of time (optimization interval) that each job $j$ spends on each resource type $i$.
  • Objective.  Given an allocation matrix $x$, we adopt a normalized effective throughput of job $j$ from POP [42], which is a linear function of $x_{*j}$. The objective is to maximize the minimum normalized effective throughput:
  • Resource constraints.  For each computing resource type $i$, the total amount of resources requested by all jobs cannot exceed the corresponding capacity:
    $\sum_{j}x_{ij}z_{j}\leq\text{capacity}_{i},\quad\forall i,$
    where $z_{j}$ is the amount of resources requested by job $j$.
  • Demand constraints.  For each job $j$, the sum of all fractions of time must not exceed 1 naturally:
    $\sum_{i}x_{ij}\leq 1,\quad\forall j.$
Proportional fairness.  The second variant of cluster scheduling maximizes the overall resource utilization while ensuring minimum service for each job. This variant shares the same variables and constraints as above, except for the objective:
Applying DeDe.  Both variants of cluster scheduling exhibit the separable structure required by DeDe. Their objectives to maximize are a sum or min of utilities for allocating each job, and a linear constraint is applied to each job and each computing resource. Thus, DeDe can decompose the problem into $n$ per-job allocations and $m$ per-demand allocations.
@@ 5.2 Traffic Engineering
Traffic engineering problems [24, 55, 44, 43, 65] allocate traffic demands between datacenters onto the network links in a wide-area network (WAN) topology.
Maximizing total flow.  This problem variant focuses on maximizing the total flow allocation.
  • Variables.  Let $x_{(u,v)(s,t)}\geq 0$ denote the amount of flow from a datacenter pair $(s,t)$ assigned to a network link $(u,v)$. In common path-based traffic engineering, flows between each node pair $(s,t)$ are allocated only over links along pre-configured paths $P_{(s,t)}$.
  • Objective.  For each node pair $(s,t)$, the inflow it incurs to node $v$ is computed as: $\text{inflow}(v,(s,t))=\sum_{(u,v)\in P_{s,t}}x_{(u,v)(s,t)}$, and outflow as: $\text{outflow}(v,(s,t))=\sum_{(v,u)\in P_{s,t}}x_{(v,u)(s,t)}$. Consequently, the total flow allocated for the node pair $(s,t)$ is captured by $\text{inflow}(t,(s,t))$, the total inflow toward the destination $t$. Therefore, the objective of maximizing the total flow between all node pairs can be expressed as:
@@
  • Resource constraints.  For each link $(u,v)$, the cumulative flow from all node pairs traversing it must not exceed its link capacity $c_{u,v}$:
@@
Minimizing max link utilization.  This variant aims to evenly distribute traffic demands across network links to avoid overloading any single link. The utilization of link is commonly defined as the (uncapped) ratio of total flow assigned to a link to the link’s capacity, although it cannot actually exceed 100% in practice. This formulation shares the same variables and constraints as above, with a revised objective:
  • Objective.  The goal is to minimize the max link utilization across the network:
@@
Applying DeDe.  The traffic engineering problem aligns with the separable structure of DeDe: the objectives are expressed as either a sum over per-demand utilities or as a maximum over per-link utilities, and linear constraints are applied to each demand and each link. Therefore, DeDe can decompose the problem into $|E|$ per-link allocations and $|V|^{2}$ per-demand allocations for parallelism. In practice, due to the large overhead of maintaining $|V|^{2}$ per-demand subproblems, we group these subproblems by their sources, reducing the total number of subproblems to just $|V|$.
@@ 5.3 Load Balancing
Load balancing problems [14, 49, 53, 44] allocate data shards among storage servers to scale out query loads in a distributed store. Load balancing is a non-convex problem with integer variables, wherein DeDe can provide a high-quality solution empirically, as discussed in §4.2.
Minimizing shard movements.  This problem aims to minimize shard movements across servers during load changes while keeping the load nearly balanced on each server.
  • Variables.  The allocation plan is encoded in a matrix $x\in[0,1]^{n\times m}$, where $x_{ij}$ is the fraction of data shard $j$ assigned to server $i$. Additionally, a binary matrix $x^{\prime}$ captures shard placement, where $x^{\prime}_{ij}=1$ if the data shard $j$ is located on server $i$ (i.e., when $x_{ij}>0$); otherwise, $x^{\prime}_{ij}=0$.
  • Objective.  The movement of data shard $j$ allocated to server $i$ can be represent as $(1-T_{ij})x^{\prime}_{ij}f_{i}$, where $T_{ij}$ denotes the initial shard placement, and $f_{j}$ denotes the memory footprint of data shard $j$. The objective of minimizing total shard movement can be expressed as:
@@
  • Resource constraints.  For each server $i$, the total query load must be close to the average query load $L$. Meanwhile, the total memory usage must be within the memory capacity:
@@
  • Demand constraints.  Each data shard $j$ must be entirely allocated across servers:
@@
Applying DeDe.  The load balancing problem also demonstrates a separable structure: the objective is the sum of allocation quality per data shard, while linear constraints are applied independently to each server and each shard. This allows DeDe to decompose the problem into $n$ per-server allocations and $m$ per-shard allocations.
@@ 6 Implementation of DeDe
Listing 1 showcases a basic resource allocation example using DeDe. Line 5 creates an $N\times M$ non-negative allocation matrix x, where x[i,j] represents the fraction of time that demand j runs on resource i. Lines 12–13 define a constraint for each resource using parameters, and Lines 14–15 define a constraint for each demand. The objective, defined in Line 18, is to maximize the sum of entries in x. Finally, the problem is constructed in Lines 21–22 and solved in Line 23.
Built on top of the popular optimization modeling language cvxpy [15, 3], DeDe inherits most of its syntax and APIs, such as Variable(), Parameter(), and Maximize(). A notable distinction is that DeDe requires users to explicitly separate resource constraints and demands constraints when initializing a problem (Line 22). As a parallel optimization tool, DeDe allows users to configure the number of CPUs (Line 23); by default, all available cores are used.
Internally, DeDe solves the optimization in three stages:
  • Problem parsing.  DeDe transforms all inequality constraints into equivalent equality constraints through non-negative slack variables. E.g., x[:,j].sum() <= 1 is converted to x[:,j].sum() + sj == 1 with a slack variable sj. Then slack variables are treated just like other variables.
  • Problem building.  DeDe organizes resource constraints into disjoint per-resource groups and demand constraints into disjoint per-demand groups. For each group, a per-resource or per-demand subproblem is constructed with cvxpy following Equations 8 and 9.
  • Problem solving.  DeDe solves these subproblems in parallel using multiple CPU cores. During ADMM iterations, only the parameters are updated, avoiding the overhead of rebuilding problems in cvxpy. Similarly, for the same problem with varying resources and demands, only the relevant parameters are updated.
1import numpy as np
2import dede as dd
3
4# Create allocation variables
5x = dd.Variable((N, M), nonneg=True)
6
7# Create parameters
8param = dd.Parameter(
9 N, value=np.random.uniform(0, 1, N))
10
11# Create constraints
12resource_constrs = [
13 x[i,:].sum() <= param[i] for i in range(N)]
14demand_constrs = [
@@
16
17# Create an objective
18obj = dd.Maximize(x.sum())
19
20# Construct and solve the problem
21prob = dd.Problem(
@@
23prob.solve(num_cpus=64, solver=dd.ECOS)
@@
We use Ray [41] to parallelize our computation across multiple cores, as Python’s native multithreading model is constrained by the (notorious) global interpreter lock (GIL). Ray circumvents this issue by managing multiple Python interpreter processes, effectively bypassing the GIL. Moreover, Ray offers robust support for inter-process communication, sparing us from manually implementing such mechanisms using Python’s lower-level multiprocessing primitives.